TOMBOW is a small manufacturer of pencils and has had the following sales record for the most recent five months: Month 1. Units sold (000) 2 3 4 250 245 260 280 320 a. Forecast demand for months 4 and 5 using a three-month moving average. b. Repeat (a) using a three-month weighted moving average, the weights are 6,3, and 1, where 6 refers to the most recent demand. al smoothing with α -0.6. Assume the forecast demand for month 3 was 247. Based on these limited data, which method do you have more confidence in? What is the forecast for month 6 using the best method? d.
Answer #1

a) 3 Month Moving Average ,

Ft = (Dt-1 + Dt-2 +Dt-3)/3

F4 = (D3 + D2 + D1)/3 = (260+245+250)/3 = 252

Period Units Sold (000) 3 MA
1 250
2 245
3 260
4 280 252
5 320 262

b ) 3 Month Weighted Average

Weighted Average Forecast formula

Ft = Wt-1*Dt-1 + Wt-2*Dt-2 + Wt-3*Dt-3

In the given Problem , Wt-1 = 0.6 , Wt-2 = 0.3 & Wt-3 = 0.1

Period Units Sold (000) 3 MWA
1 250
2 245
3 260
4 280 255
5 320 271

3.

alpha,\alpha = 0.35

Formula for calculating the forecast using exponential smoothing method

Ft+1 = \alpha *At + (1-\alpha)*Ft

Where Ft+1 is forecast value for t+1 period

At = Actual demand for t period

Ft is forecast value for t period

Period Units Sold (000) ES(0.6)
1 250
2 245
3 260 247
4 280 255
5 320 270

d) Calculate MAD for each technique

MAD = Sum of Absolute Error/ Number of Period

Error = Ft - Dt

Absolute Error = Absolute value of (Ft - Dt )

Find Error and Absolute error of each Period for all the forecasting type

Period Units Sold (000) 3 MA Absolute(Actual-Forecast)
1 250
2 245
3 260
4 280 252 28.3
5 320 262 58.3
MAD 43.3
Period Units Sold (000) 3 MWA Absolute(Actual-Forecast)
1 250
2 245
3 260
4 280 255 25.5
5 320 271 49.5
MAD 37.5
Period Units Sold (000) ES(0.6) Absolute(Actual-Forecast)
1 250
2 245
3 260 247
4 280 255 25.2
5 320 270 50.08
MAD 37.6

Technique with less MAD is better technique . So 3 MWA technique can give more confidence

Period Units Sold (000) 3 MWA
1 250
2 245
3 260
4 280 255
5 320 271
6 302

F6 = 302